Good Evening. Please help me solve this system of linear equation:
(3/2)x+(2/3)y=1
(2/3)x-(3/2)y=0
Thnak You!

Answer by Boreal(15235)

(3/2)x+(2/3)y=1
(2/3)x-(3/2)y=0
multiply both equations by 6
9x+4y=6
4x-9y=0
top by 9, bottom by 4
81x+36y=54
16x-36y=0
97x=54
x=54/97
-------------
top by -4, bottom by 9
-36x-16y=-24
36x-81y=0
-97y=-24
y=24/97
-------------------
check
486/97+96/97=582/97=6
216/97-216/97=0